The quarterly sales data (number of book sold) for Christian book over the past three years in California follow: (You can use Excel to compute the equation)

Quarter

Year 1

Year 2

Year 3

1

2

3

4

1230

1020

2534

2600

1470

990

2800

2590

1520

1020

2850

2700

1. Use the following dummy variables to develop an estimated regression equation to account for any seasonal effects in the data: Quarter1=1 if the sales data point is in Quarter 1, otherwise Quarter 1=0; Quarter 2=1 if the sales data point is in Quarter 2, otherwise, Quarter 2=0; Quarter 3=1 if the sales data point is in Quarter 3, otherwise Quarter 3=0.

2. Compute the quarterly forecasts for next year.

3. Let t=1 to refer to the observation in quarter 1 of year 1; t=2 to refer to the observation in quarter 2 of year 1;,,,, and t=12 to refer to the observation in quarter 4 of year 3. Using the dummy variables defined in part (2) and t, develop an estimated regression equation to account for seasonable effects and any linear trend in the time series. Based upon the seasonal effects in the data and linear trend, compute the quarterly forecasts for next year.
Answer #1

Period(t)

Actual Data(y)

Y*t

t2

1

1230

1230

1

2

1020

2040

4

3

2534

7602

9

4

2600

10400

16

5

1470

7350

25

6

990

5940

36

7

2800

19600

49

8

2590

20720

64

9

1520

13680

81

10

1020

10200

100

11

2850

31350

121

12

2700

32400

144

Total

78

23324

162512

650

t-bar = sum(t)/n = 78/12 = 6.5

y-bar = sum(y)/n = 23324/12= 1943.67(Rounding to 2 decimal places)

Regression equation is y = a + bt where a = intercept and b = slope

b = (sum (y*t) – n* t-bar *y-bar)/(sum(t2) – n*t-bar2)

= (162512–12*6.5*1943.67)/(650 – 12*6.5*6.5) = (162512-151606.26)/(650 – 507) = 10905.74/143= 76.26391608 = 76.26 (Rounding to 2 decimal places)

a = y-bar – b*x-bar = 1943.67 – 76.26*6.5 = 1447.98

So, regression equation becomes y = 1447.98 + 76.26t

Plotting t = 1,2….20 we get de-seasonalised values at ydt

So the table becomes

Period(t)

Actual Data(y)

Yt

t^2

Deseasonalized data(ydt)

Seasonality factor

1

1230

1230

1

1524.24

0.806959534

2

1020

2040

4

1600.5

0.637300843

3

2534

7602

9

1676.76

1.511247883

4

2600

10400

16

1753.02

1.483154784

5

1470

7350

25

1829.28

0.803594857

6

990

5940

36

1905.54

0.519537769

7

2800

19600

49

1981.8

1.412856999

8

2590

20720

64

2058.06

1.258466711

9

1520

13680

81

2134.32

0.712170621

10

1020

10200

100

2210.58

0.461417366

11

2850

31350

121

2286.84

1.246261216

12

2700

32400

144

2363.1

1.142566967

Total

78

23324

162512

650

  1. Seasonality = Actual data /deseasonalized data

Example Seasonality for period 1 = 1230/1524.24=0.806959534

Using periodicity = 4 we can say seasonal factor of Quarter 1 would be average of seasonal factors of period 1,5,and 9 = Average(0.806959534, 0.803594857, 0.712170621)

= 0.774241671

Likewise seasonality factors for all quarters are calculated

Quarter

Seasonal factor

1

0.774241671

2

0.539418659

3

1.390122033

4

1.294729488

  1. Quarterly forecast for next year

Quarter t = (1447.98 + 76.26t)* seasonality for quarter 1 [ Next year Quarter 1 would be period 13 so value of t = 13 and seasonality for quarter 1 would be applied here due to periodicity]

So,

Next year-Quarter 1 = (1447.98 + 76.26*1)* 0.774241671= 1180.130125

Next year-Quarter 2 = (1447.98 + 76.26*2)* 0.539418659= 863.3395637

Next year-Quarter 3 = (1447.98 + 76.26*3)* 1.390122033= 2330.901

Next year-Quarter 4 = (1447.98 + 76.26*4)* 1.294729488= 2269.687